In a quiz, positive marks are given for correct answers and negative marks are given for incorrect answers. If Jack’s scores in five successive rounds were 25, -5, -10, 15 and 10, what was his total at the end?

Grade:12th pass

1 Answers

Pawan Prajapati
2 years ago
Given scores are 25, -5, -10, 15, 10 Marks given for correct answers = 25 + 15 + 10 = 50 Marks given for incorrect answers = (-5) + (-10) = -15 ∴ Total marks given at the end = 50 + (-15) = 50 – 15 = 35